# CR-Net: Scaling Parameter-Efficient Training with Cross-Layer Low-Rank Structure

### v1

@@ A.2 Computation analysis without re-computation
Appendix B Computation and memory analysis with re-computation
B.1 Activation memory and re-computation analysis of CR-Net
@@ B.1.2 Computation analysis
B.2 Re-computation complexity of CoLA-M
Table 2 presents the recomputation complexity of activation-effective approaches, including CoLA-M. In fact, the re-computation complexity of CoLA-M can be obtained by the similar analysis process of CR-Net in Appendix B.1.2. The total re-computation complexity for one gradient step of CoLA-M is:
@@ B.2 Re-computation complexity of CoLA-M
Appendix C Theoretical insight: Residual Subtraction Improves Low-Rank Approximation
@@
Appendix D Experimental details and additional experiments
D.1 Experimental details for the observation of cross-layer low-rank structure
@@
We subsequently quantified the activation approximation errors through two distinct methodologies: (a) direct low-rank estimation $\text{LR}_{r}(Y_{l}^{\text{P}})$ and (b) matrix recovery via Equation (3). The rank parameter $r=0.25h,0.5h$ respectively. And for $\widetilde{Y}_{l,\beta_{0}}^{\text{P}}$ definited in Eq. (3), the scaling factor $\beta_{0}$ is tuned from $\{0,0.2,0.4,0.6,0.8,1,2,3,5\}$. And the other hyperparameters can be referred from Table 6.
@@
D.2 Experimental details for C4 pre-training tasks
D.2.1 Experimental set up
@@
D.2.2 CR-Net accelerates the pre-training tasks
@@
Table 9: Comparsion of validation perplexity ($\downarrow$) of different approaches in LLaMA-2 1B pre-training tasks. The results of compared methods are referred from [42, 66].
@@
Table 10: Comparsion of validation perplexity ($\downarrow$) of different approaches in LLaMA-2 13B pre-training tasks.
@@
Table 11: Comparsion of validation perplexity ($\downarrow$) and accuracy ($\uparrow$) of fine-tuning tasks with different approaches.
@@
D.2.3 Scaling CR-Net to a larger model size
To validate the performance of CR-Net with a larger model size, we pre-train a LLaMA-2 13B model under 8-bit Adam and CR-Net separately. We use the re-computation present in Section 3.3 for CR-Net and save a series of full activation every 8 layers. We trained 40K steps due to the time limitation. The hyperparameter rr in CR-Net is set to 1260. The evaluation perplexity shown as Table 10 illustrates the performance of CR-Net in 13B model, which demonstrates that CR-Net exhibits a comparable performance than full-rank training with 8-bit Adam. Thus can be scaled to models with 13B parameters.
D.2.4 Downstream performance
@@
D.2.5 Experimental setting and additional results for the ablations
@@
How does rank selection impact pre-training performance? In this ablation, we train LLaMA-2 350M with CR-Net for 60K iterations. The max learning rate is set to $0.01$ and all the other configs is the same as that in Appendix D.2.1. For the selection of layer ranks, we present the following three strategies to ensure the their parameters are all the same:
@@
Figure 9: The comparsion of evaluation perplexity for CR-Net in pre-training tasks with fixed $\beta_{l}^{\text{P}}$ and learnable $\beta_{l}^{\text{P}}$. (Left: LLaMA-2 350M, right: LLaMA-2 1B.)
Whether other cross-layer residual strategies do well in pre-training with low-rank parameters? To compare CR-Net with different efficient cross-layer residual strategies for LLMs pre-training including ResFormer [65] and DenseFormer [49], we pre-trained a LLaMA-2 1B model with CR-Net, Learnable-ResFormer, and DenseFormer using low-rank parameters. The rank $r$ for transformer layers (except the first layer) is set to 448, while all other configs are the same as those in Appendix D.2.1. We stopped training early at the 50,000-th step.
@@
Table 12: Comparsion of validation perplexity ($\downarrow$) of different approaches in LLaMA-2 pre-training tasks with longer sequence length and more training tokens.
@@
D.2.6 Pre-training with longer sequence and more training tokens
@@
Table 14: Comparsion of validation perplexity ($\downarrow$) of different approaches in LLaMA-3 pre-training tasks.
@@
Table 15: Comparsion of validation perplexity ($\downarrow$) of different approaches in Qwen-3-based pre-training tasks.
@@
D.3 Pre-training GQA and MoE models with CR-Net
@@
Pre-training LLaMA-3 with GQA architecture. We pre-trained a 1B-parameter LLaMA-3 model on 15.6B tokens from the C4-en dataset using both GaLore and CR-Net methods. The model hyperparameters are shown in Table 13. The low-rank coefficient rr was set to 512 for GaLore and 448 for CR-Net. The remaining configuration follows the same setup as the LLaMA-2 pre-training task described in Appendix D.2. The evaluation perplexity results are presented in Table 14, indicating that CR-Net outperforms GaLore in pre-training tasks on the LLaMA-3 model.
Pre-training models with MoE architecture. We pre-trained a Qwen-3-based model under the MoE architecture, containing 1.8B total parameters with 650M activated parameters, on 7.8B tokens from the C4-en dataset. Training was conducted under both full-rank and CR-Net settings. The model hyperparameters are provided in Table 5. The low-rank coefficient rr in CR-Net was set to 224. For CR-Net, we replaced the parameter matrices within the linear operators of the MoE layers. Each MoE layer was treated as a unified operator, and cross-layer residual connections were applied after obtaining the output of the MoE operators. The remaining experimental setup matches that of the LLaMA-2 pre-training task in Appendix D.2. The evaluation perplexity, shown in Table 15, demonstrates the competitive performance of CR-Net compared to full-rank training.
Appendix E Additional discussions
E.1 Per-Step Computation Acceleration Analysis.
@@ Appendix E Additional discussions
E.2 Additional discussion for the learnable scaling factor $\beta_{l}^{\text{P}}$.
@@
Appendix F Multi-GPU training: a theoretical analysis for CR-Net with pipeline parallelism
While the single-GPU throughput results (Figure 6) validate the computational efficiency of CR-Net , multi-GPU training—especially with parallelism strategies like Pipeline Parallelism (PP) [46, 47, 28, 54]—introduces critical considerations around communication overhead. Cross-layer residual connections, a core component of CR-Net, could theoretically increase data transmission between GPUs, as forward/backward passes require sharing activation residuals across worker nodes. To address this concern, we conduct a quantitative analysis of computation-communication tradeoffs for CR-Net in PP-enabled multi-GPU setups. We also present an analysis of HBM memory with pipeline parallelism.

### v2

@@ B.2 Computation analysis without re-computation
Appendix C Computation and memory analysis with re-computation
C.1 Activation memory and re-computation analysis of CR-Net
@@ C.1.2 Computation analysis
C.2 Re-computation complexity of CoLA-M
Table 2 presents the recomputation complexity of activation-effective approaches, including CoLA-M. In fact, the re-computation complexity of CoLA-M can be obtained by the similar analysis process of CR-Net in Appendix C.1.2. The total re-computation complexity for one gradient step of CoLA-M is:
@@ C.2 Re-computation complexity of CoLA-M
Appendix D Theoretical insight: Residual Subtraction Improves Low-Rank Approximation
@@
Appendix E Experimental details and additional experiments
E.1 Experimental details for the observation of cross-layer low-rank structure
@@
We subsequently quantified the activation approximation errors through two distinct methodologies: (a) direct low-rank estimation $\text{LR}_{r}(Y_{l}^{\text{P}})$ and (b) matrix recovery via Equation (3). The rank parameter $r=0.25h,0.5h$ respectively. For $\widetilde{Y}_{l,\beta_{0}}^{\text{P}}$ defined in Eq. (3), the scaling factor $\beta_{0}$ is selected from the set $\{0,0.2,0.4,0.6,0.8,1,2,3,5\}$ to minimize the relative error given by (4). And the other hyperparameters can be referred from Table 8.
@@
E.2 Experimental details for C4 pre-training tasks
E.2.1 Basic experimental set up
@@
Experimental setup for training and inference throughput evaluation. To measure throughput performance, we pre-trained the LLaMA-2 1B model on an A100 40G GPU with a batch size of 16. For CR-Net , we evaluated both single-GPU configurations and multi-GPU scenarios employing data parallelism across 4 GPUs. For inference throughput measurements, we utilized a A100 80G GPU with a microbatch size of 64. Additionally, when assessing throughput for the LLaMA-2 7B model, we employed a A100 80G GPU with a batch size set to 512. In all experiments, we utilized the maximum microbatch size that could be accommodated without triggering out-of-memory errors.
E.2.2 CR-Net accelerates the pre-training tasks
@@
Table 11: Comparison of validation perplexity ($\downarrow$) of different approaches in LLaMA-2 1B pre-training tasks. The results of compared methods are referred from [43, 69].
@@
Table 12: Comparison of validation perplexity ($\downarrow$) of different approaches in LLaMA-2 13B pre-training tasks.
@@
Table 13: Comparison of validation perplexity ($\downarrow$) for CR-Net with and without activation quantization in LLaMA-2 350M pre-training tasks.
Training tokens 6.4B
Full-rank 18.80
CR-Net 18.95
CR-Net with activation quantization 19.32
Table 14: Comparison of validation perplexity ($\downarrow$) and accuracy ($\uparrow$) of fine-tuning tasks with different approaches.
@@
E.2.3 Scaling CR-Net to a larger model size
To validate the performance of CR-Net with a larger model size, we pre-train a LLaMA-2 13B model under 8-bit Adam and CR-Net separately. We use the re-computation present in Section 3.3 for CR-Net and save a series of full activation every 8 layers. We trained 40K steps due to the time limitation. The hyperparameter rr in CR-Net is set to 1260.
The evaluation perplexity shown as Table 12 illustrates the performance of CR-Net in 13B model, which demonstrates that CR-Net achieves more than 50% reduction in parameters, while incurring only a 2% degradation in validation performance.
E.2.4 Pre-training CR-Net with activation quantization
To validate the performance of CR-Net with activation quantization, we pre-train a LLaMA-2 350M model under activation quantization. We adopt the same quantization scheme as Q-LoRA [15], which employs NF4 precision and the DoubleQuant strategy with a block size of 16. Table 13 presents the evaluation perplexity of CR-Net with and without activation quantization, compared to full-rank training. The results indicate that CR-Net with 4-bit activation quantization achieves additional activation memory savings while incurring less than a 3% increase in perplexity.
E.2.5 Downstream performance
@@
E.2.6 Experimental setting and additional results for the ablations
@@
How does rank selection impact pre-training performance? In this ablation, we train LLaMA-2 350M with CR-Net for 60K iterations. The max learning rate is set to $0.01$ and all the other configs is the same as that in Appendix E.2.1. For the selection of layer ranks, we present the following three strategies to ensure the their parameters are all the same:
@@
Figure 9: The Comparison of evaluation perplexity for CR-Net in pre-training tasks with fixed $\beta_{l}^{\text{P}}$ and learnable $\beta_{l}^{\text{P}}$. (Left: LLaMA-2 350M, right: LLaMA-2 1B.)
Whether other cross-layer residual strategies do well in pre-training with low-rank parameters? To compare CR-Net with different efficient cross-layer residual strategies for LLMs pre-training including ResFormer [68] and DenseFormer [52], we pre-trained a LLaMA-2 1B model with CR-Net, Learnable-ResFormer, and DenseFormer using low-rank parameters. The rank $r$ for transformer layers (except the first layer) is set to 448, while all other configs are the same as those in Appendix E.2.1. We stopped training early at the 50,000-th step.
@@
Table 15: Comparison of validation perplexity ($\downarrow$) of different approaches in LLaMA-2 pre-training tasks with longer sequence length and more training tokens.
@@
E.2.7 Pre-training with longer sequence and more training tokens
@@
Table 17: Comparison of validation perplexity ($\downarrow$) of different approaches in LLaMA-3 pre-training tasks.
@@
Table 18: Comparison of validation perplexity ($\downarrow$) of different approaches in Qwen-3-based pre-training tasks.
@@
E.3 Pre-training GQA and MoE models with CR-Net
@@
Pre-training LLaMA-3 with GQA architecture. We pre-trained a 1B-parameter LLaMA-3 model on 15.6B tokens from the C4-en dataset using both GaLore and CR-Net methods. The model hyperparameters are shown in Table 16. The low-rank coefficient rr was set to 512 for GaLore and 448 for CR-Net. The remaining configuration follows the same setup as the LLaMA-2 pre-training task described in Appendix E.2. The evaluation perplexity results are presented in Table 17, indicating that CR-Net outperforms GaLore in pre-training tasks on the LLaMA-3 model.
Pre-training models with MoE architecture. We pre-trained a Qwen-3-based model under the MoE architecture, containing 1.8B total parameters with 650M activated parameters, on 7.8B tokens from the C4-en dataset. Training was conducted under both full-rank and CR-Net settings. The model hyperparameters are provided in Table 7. The low-rank coefficient rr in CR-Net was set to 224. For CR-Net, we replaced the parameter matrices within the linear operators of the MoE layers. Each MoE layer was treated as a unified operator, and cross-layer residual connections were applied after obtaining the output of the MoE operators. The remaining experimental setup matches that of the LLaMA-2 pre-training task in Appendix E.2. The evaluation perplexity, shown in Table 18, demonstrates the competitive performance of CR-Net compared to full-rank training.
Appendix F Additional discussions
F.1 Per-Step Computation Acceleration Analysis.
@@ Appendix F Additional discussions
F.2 Additional discussion for the learnable scaling factor $\beta_{l}^{\text{P}}$.
@@
Appendix G Multi-GPU training: an analysis for CR-Net with pipeline parallelism
While the single-GPU throughput results (Figure 3) validate the computational efficiency of CR-Net , multi-GPU training—especially with parallelism strategies like Pipeline Parallelism (PP) [49, 50, 29, 57]—introduces critical considerations around communication overhead. Cross-layer residual connections, a core component of CR-Net, could theoretically increase data transmission between GPUs, as forward/backward passes require sharing activation residuals across worker nodes. To address this concern, we conduct a quantitative analysis of computation-communication tradeoffs for CR-Net in PP-enabled multi-GPU setups as well as an empirical verification. We also present an analysis of HBM memory with pipeline parallelism.
@@ Appendix G Multi-GPU training: an analysis for CR-Net with pipeline parallelism
Empirical verification. To empirically validate communication efficiency, we pre-trained a LLaMA-2 70B model with CR-Net under pipeline parallelism using 8 NVIDIA A800 GPUs interconnected via NVLink, on the C4-en dataset (sequence length=256). Pipeline parallel training was implemented with the torch.distributed.pipelining library, with each GPU handling 10 transformer layers.
The measured communication time for one forward and backward step was 1.380 seconds, yielding a bandwidth of 192.5 GB/s. This bandwidth is approximately three times higher than the PCIe bandwidth assumed in our theoretical analysis (Table 22). These findings demonstrate that under NVLink connectivity, CR-Net achieves shorter communication times than theoretically projected, further confirming that for large scale models, the communication volume does not substantially impair overall efficiency, thus supporting our theoretical conclusions.